OpenStudy (anonymous):

Question: Use the formulas for lowering powers to rewrite the expression in terms of the first power of cosine cos4 x sin2 x Answer: 1/16x - 1/64sin(4x) + 1/48 sin^3(2x) + C Why is this answer wrong in my math software that I am using for school. I tried integrating the problem. Can anyone help by explaining step by step if it's possible, Please thank you?

OpenStudy (anonymous):

i'm assuming you mean cos^4x sin^2x

OpenStudy (anonymous):

cos^2x=(1+cos2x)/2

OpenStudy (anonymous):

sin^2x=(1-cos2x)/2

OpenStudy (anonymous):

((1-cos2x)/2)((1+cos2x)/2)^2

OpenStudy (anonymous):

expand

OpenStudy (anonymous):

1/8(1+cosx+cos^2(2x)+cos^3(2x))

OpenStudy (anonymous):

split up the functions.

OpenStudy (anonymous):

into \[1/8(\int\limits(1+\cos x+\cos ^{2}2x)dx+\int\limits( \cos ^{2}2x*\cos2x)dx\]

OpenStudy (anonymous):

then u substitution.

OpenStudy (anonymous):

u=sin2x

OpenStudy (anonymous):

then it's easy after that.

OpenStudy (anonymous):

i got lost after the 8th statement